Find the magnitude of the vector. (-7, -3)
Answer #1

The magnitude of a vector, in the form of (x,y) is given by the formula:
Magnitude of vector = \sqrt{x^2+y^2}

Here, x=-7,y=-3

Therefore, the magnitude of the vector will be

\sqrt{(-7)^2+(-3)^2}=\sqrt{49+9}=\sqrt{58}

Hence, the magnitude of the vector (-7,-3) will be \sqrt{58}